OpenStudy (anonymous):

-4x^3 -24 -20 Factor the following expression I've gotten the x+2 and x+3

OpenStudy (anonymous):

well first you would set it equal to 0 (-4x^3-24-20=0) then you would subtract -20 from -24 and get -44 (-4x^3-44=0) then factor out 4 ( 4(x^3-11=0) ) then solve x^3-11=0 by subtracting 11 to get (x^3= 11) then cube 11 and get \[+-\sqrt{11}\]

OpenStudy (anonymous):

put that in your calculator

OpenStudy (anonymous):

...

OpenStudy (leilanilane):

Okay The first thing I notice is that it has factors Yay!! Okay so this is what we do: -4x^3 - 24 - 20 -4(x^3 + 6 + 5) Should be the correct answer Unless your trying to split the middle part of the expression into two parts

OpenStudy (leilanilane):

lalaz888 is all wrong lol